# Frequency Is What You Need: Word-frequency Masking Benefits Vision-Language Model Pre-training

### v2

@@ 5 The Importance of Frequency Distributions
Figure 3: The figure illustrates the distribution of top-50 words in the text before and after applying various text masking strategies. We set the text length after text masking to 6. The x-axis represents the word index, which is sorted by counts of the original data, and the y-axis shows the word frequency. The dataset used is CC12M and the value of $t$ of formula 2 is set to $10^{-6}$. We remove special characters from the vocabulary.
Next, we look at how the different masking strategies impact the frequency distribution of words in the training data. Fig. 3 shows the distribution of words in the text before and after applying all of the text masking strategies. We see that in general, there is quite a large difference between the frequency of the highest-frequency words and the frequency of words with lower frequency. Such a long-tail distribution is characteristic of words occurring in text, which is known to be Zipf distributed. Most masking approaches preserve the general shape of this distribution: The frequency of a word in the masked training data tracks its frequency in the original data. The exception is syntax masking, which shows peaks and pits for specific words. Specifically, because syntax masking retains more nouns, the articles “the” and “a” are less frequent, but the nouns create spikes. The rationale for prioritizing nouns is solid. Nouns are closely related to the depicted content of images and, for this reason, are considered to be more helpful. Indeed, the helpfulness of nouns for the VLM is reflected in the fact that syntax masking learns faster than other text masking strategies. In Fig. 1, we saw that syntax masking outperforms the other masking approaches during the early epochs. However, retaining too many nouns can have two adverse effects. First, the model can overfit to certain high frequency nouns. In Fig. 3, we see that CLIPA syntax masking retains more occurrences of “person,” “illustration,” and “stock” than the other strategies. Second, retaining nouns means necessarily dropping words from other parts-of-speech (POS) categories, which may be important, as discussed in more detail below.
Considering the CLIPA truncation masking strategy, we see in Fig. 1 that the word frequency distribution of the truncated training data shows a difference with the other masking strategies. Specifically, truncation is the top line on the left of the graph, which means truncation is making the most highly frequent words in the training data relatively even more frequent. The fact that truncation impacts word frequency might seem surprising at first, but it can be explained if we consider that the probability of a word occurring changes with sentence position. With truncation, typical text beginnings, e.g., “A person is…”. remain over-represented in the data.
To dive deeper into the importance of the distribution of words over POS categories, we provide statistics for each of the different text masking strategies in Table 5. Consulting the column “Other”, we see an interesting pattern. Syntax masking changes the proportion of words in the “Other” category radically with respect to the data before masking, but truncation, random and block do not. CLIPF appears to provide a happy medium. CLIPF retains more nouns than CLIPA truncation, random, and block, but not so many that the “Other” category becomes poorly represented.
Table 5: To examine the distribution of words belonging to different POS categories, we calculate the percentage of words in the training data belonging to each category in before and after applying text masking. The dataset is CC12M, and we retain 6 words for each text after applying text masking.
@@
The importance of “Other” words is surprising because the category includes many words, which like articles, lack semantic content and are sometimes considered to have no direct connection to images. In fact, in many NLP applications, articles and other “stop words” are simply removed. However, since the CLIP text encoder is a transformer, technically it should be able to exploit the context of content words, particularly words in “Other” category. Note that with syntax masking a phrase like “smoking in the room” in the training data would always be reduced to “smoking room”, which has a different meaning. However, with the other masking approaches, “in the” would not be consistently deleted. Note specifically that block masking retains entire phrases, without always using the beginning of the sentence as truncation masking does. Also, frequency masking (CLIPF) retains entire phrases in some cases, more than would be expected with random masking.
In sum, when trained with enough epochs, overfitting effects with syntax masking occur. Our analysis has related this problem to peaks in the word frequency distribution caused by over-retention of nouns. Also, with enough epochs, truncation masking overfitting effects occur, which our analysis has related to distortion of the word frequency distribution caused by oversampling of the beginning of sentences. In sum, our analysis of word frequency distributions in the training data provides insight into why the performance of syntax and block masking falls off in Fig. 1. Further, if word frequency is indeed important, then masking that controls word-frequency directly, such as our CLIPF, can be considered to be a sensible and transparent approach.
6 Experimental Results
In this section, we carry out a systematic investigation of the existing text masking approaches from CLIPA [12] and our CLIPF approach. Our results support our observation that syntax and truncation masking are not productive approaches to masking and highlight the benefits of masking that directly leverages frequency information, i.e., our CLIPF. Our experiments use 30 epochs, following Open_CLIP [9]. We note that VLM training work has shown performance improvements up to 100 epochs [5, 11] and we would expect that with more epochs the performance would also improve. However, the improvement beyond 30 is expected to be small, predictable and not necessarily important in an operational setting where conserving compute is of concern.
@@ 6.1 Zero-shot Performance on ImageNet-1k
Table 4 presents results for zero-shot classification on ImageNet-1K for models trained on both the CC3M and CC12M data. The experiments compare different masking approaches and different masking ratios, indicated in the table with the number of input text tokens (Text Tokens column). Models are pre-trained for 30 epochs (pre-train column) and then fine-tuned for one epoch (fine-tune column) with the full, unmasked data. Note that since the average text length in CC3M is less than 16 tokens, the benefits of fine-tuning a model pre-trained with 16 text tokens are limited.
We see that CLIPF outperforms the other CLIPA matching approaches in nearly all cases across text masking ratios (input sizes) as reflected by the bolded accuracy scores. In general, random, syntax, and block strategies exhibit similar performance; however, truncation performs worse. Contrary to the conclusion of the original CLIPA work [12], syntax masking does not yield the best performance.
It is particularly interesting to observe from the results in Table 4 the fact that models trained with text masking (CLIPA and CLIPF) are competitive with, or superior to, models trained on the original, unmasked text data (CLIP, FLIP). Note that at a text masking ratio of 75% (corresponding to 8 tokens), the masking approaches are 1.3 $\times$ faster than FLIP and 4.0 $\times$ faster than CLIP.
We first consider the comparison with CLIP. On CC3M, CLIPF approaches the performance of the original CLIP approach with a 81% text masking ratio (4 input tokens). On CC12M, CLIPF outperforms the original CLIP (no masking) with a 75% masking ratio (8 input tokens). Fine-tuned CLIPF with a 75% masking ratio achieves an accuracy of 39.3% compared to 36.6% for the original CLIP trained on the full data.
Next, we consider the comparison with FLIP, which is the same as CLIP except that it uses a 75% image masking ratio (49 image tokens). On CC3M, fine-tuned CLIPF is competitive with or outperforms fine-tuned FLIP at all masking ratios. On CC12M, fine-tuned CLIPF outperforms fine-tuned FLIP, except for the highest masking ratio (87.5% corresponding to 4 input tokens). Additionally, as shown in the last row of Table 4, if we double the number of image tokens, then CLIPF substantially beats the original CLIP (41.0% vs 36.6%). However, it also notes that more resources are required (Tokens processed from 25% to 46%).
@@
Finally, we consider the pattern of the text masking approaches as the masking ratio increases, i.e., as we decrease the input size from 16 to 4 tokens. Here, we focus on the results of the fine-tuned model. Moving from 16 to 8 input tokens benefits most approaches. Recall that the reduction of the number of input tokens allows for an increase in the batch size. As we reduce the number of tokens from 8 to 6 to 4, syntax masking and truncation do not perform as well as the other text masking strategies. We see that compared to existing text matching approaches, CLIPF is better positioned to deliver benefits as the number of input tokens decreases.
@@ 6.2 Zero-shot Robustness Evaluation
Following the evaluation of CLIP and FLIP [3, 8], we test the masking strategies across data sets. Table 6 presents results for zero-shot classification for six commonly used data sets. The performance of these strategies on multiple datasets is similar to the behavior observed on the ImageNet-1K dataset. Across different text masking ratios and across different data sets, CLIPF outperforms the other masking strategies in nearly all cases. Reducing the text tokens in CLIPF from 16 to 8 results in substantial performance improvements compared to the other text masking approaches across almost all datasets: ImageNet-A [31] by 1.28%, ImageNet-O [32] by 0.80%, ImageNet-R[33] by 4.9% , ImageNet-Sketch [34] by 3.14%, ImageNet-V2 [35] by 3.38%, and ObjectNet [36] by 0.44%. Again, we see that the benefits of CLIPF are particularly evident as the masking number of input tokens decreases.
We also observe that CLIPF is able to improve over the FLIP baseline without masking in a majority of cases. However, despite being the strongest text-masking strategy, the results reveal that it is challenging for CLIPF to improve over the original CLIP. In only a handful of cases does it outperform the original CLIP, but might do so with more image tokens as we saw in 6.1.
Table 7: Zero-shot Image-Text Retrieval. We evaluated CLIP, FLIP, CLIPA, and CLIPF image-text retrieval performance on the COCO and Flickr30k datasets. The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
Models Masking Text Tokens Text Retrieval Image Retrieval
@@
CLIP 32 62.62 35.54 45.42 24.22
FLIP 32 54.73 29.34 38.62 20.88
CLIPA truncation 16 44.67 25.54 34.99 19.64
random 16 58.48 32.36 43.61 22.81
block 16 56.51 30.82 44.06 22.66
syntax 16 54.54 29.60 41.16 21.40
CLIPF frequency 16 57.89 31.52 42.72 22.57
CLIPA truncation 8 30.47 16.92 23.96 12.78
random 8 58.58 36.24 43.79 23.16
block 8 60.06 35.88 45.98 24.65
syntax 8 50.30 29.64 38.46 20.28
CLIPF frequency 8 58.68 34.74 44.58 23.31
CLIPA truncation 6 32.25 16.18 22.39 11.93
random 6 55.23 32.58 42.47 21.34
block 6 54.44 33.12 41.68 21.73
syntax 6 46.15 26.78 34.08 17.99
CLIPF frequency 6 56.02 32.32 41.05 21.28
CLIPA truncation 4 30.97 17.44 21.09 11.07
random 4 41.62 24.14 30.26 15.63
block 4 40.83 23.56 30.77 14.98
syntax 4 38.66 21.68 26.08 14.14
CLIPF frequency 4 45.07 25.60 31.72 15.96
6.3 Zero-shot Image-Text Retrieval
Table 7 shows image/text retrieval results on Flickr30k [37] and COCO [38]. In general, we see that text masking approaches are able to improve over FLIP. When using 16 text tokens, all strategies except the truncation method outperform FLIP, which is pre-trained on the full text. Here, we do not see consistent improvement over CLIP, although some cases come lose. Note again that we are testing here only with 49 image tokens and 98 would fare better, with additional computational cost.
We point out that for Image-Text Retrieval the performance of syntax masking is particularly low in comparison to the other forms of masking. Specifically, when using 8 and 6 text tokens, the performance gap between syntax masking and block, random, and frequency masking is approximately 9% on COCO and 6% on Flickr30k for image-text retrieval. This gap is wider than what was observed with Zero-shot Classification. We conjecture that the bad performance is due to the Image-Text Retrieval task requiring understanding the overall meaning and logic of the text rather than just recognizing nouns or objects. For this reason, the fact that syntax masking retains a high number of nouns, as shown in Section 5, is particularly harmful.
@@ 6.3 Zero-shot Image-Text Retrieval
6.4 Zero-shot Classification on More Datasets
In Table 8, we present evaluation results on additional datasets. With six text tokens, CLIPF consistently outperforms the original CLIP (no masking), FLIP (image masking with full-text), and other text masking strategies. Aligned with ImageNet findings, reducing text tokens from 16 to 8 significantly boosts performance. Improvements include SUN397 [39] (+6.39%), DTD [40] (+9.31%), OxfordPets [41] (+4.64%), STL10 [42] (+9.31%), EuroSAT [43] (+6.60%), UCF101 [44] (+6.68%). The performance of other text masking strategies, including truncation, random, block, and syntax, also improves with this reduction. However, their improvements are more limited.
We point out that the Describable Textures Dataset (DTD)[40] contains 47 categories of texture patches, whose POS is mostly verb and that the OxfordPets dataset[41] also contains many class names that are not nouns. On these data sets the performance of syntax masking is particularly low. Specifically, when using six text tokens, syntax masking underperforms CLIPF by 14.10% on DTD and 21.04% on OxfordPets. This low performance reflects that the VLM has overfitted on the large number of nouns retained by syntax masking.
6.5 Scaling to the larger Dataset.
Finally, we carry out an exploratory experiment to gain insight into different masking strategies with data sets of different sizes. We carry out the largest experiment that is possible given our computational capacity, which was training on the LAION-400M [17] data set, using 4 text tokens, a smaller image size (112x112) and 50% image masking. We trained for 16 epochs. Previous works have shown that both smaller image sizes and image masking are effective for accelerating CLIP pre-training [21, 8, 45]. Results are shown in Fig. 4 for both top-1 accuracy (left) and top-5 accuracy (right). We see from these graphs that using $<$ 10 epochs would not provide a complete picture on how the different masking approaches perform relative to each other. Overall, CLIPF, the frequency based approach dominates. However, syntax is performing better than it did on smaller data sets, possibly because the large amount of data is compensating for the overemphasis of syntax masking on nouns. It is important not to conclude on the basis of this experiment that syntax masking will continue to dominate block masking. Looking at the top-5 performance, we see that these two approaches near each other already in the 16 epochs we are able to compute. In sum, this experiment shows that CLIPF is a good performer as data scales up, but that the number of epochs is important in understanding the strength of different masking strategies.
Figure 4: Zero-shot classification accuracy on ImageNet-1K over training epochs for CLIPF and CLIPA strategies when pre-training on LAION-400M datasets. We pre-trained the model using 25 image tokens and 4 text tokens for 16 epochs.

### v3

@@ 3.2 Impact of masking on word frequency
Figure 3: The figure illustrates the distribution of top-25 words in the text before and after applying various text masking strategies. We set the text length after text masking to 6. The x-axis represents the word index, which is sorted by counts of the original data, and the y-axis shows the word frequency. The dataset used is CC12M and the value of $t$ of Eq. 3 is set to $10^{-6}$. We remove special characters from the vocabulary. The figure with more words is provided in the supplementary material.
Existing literature reports the impact of text masking in terms of the percentage of words or tokens retained in the training data. Seen from the word-count perspective, the impact is the same: for any text masking approach, the proportion of words removed controls the computational savings. However, to determine the reasons why some text masking approaches are better than others, it is necessary to look beyond overall word count.
In this section, we carry out an analysis that goes beyond word count and considers the impact of text masking on the frequency of words occurring in the training set. We plot word-frequency curves, used in linguistics, for the four existing approaches to text masking and display the head of the curve (the first 25 most frequent words) in Fig. 6. The curves generally follow the long-tail distribution expected of natural language and predicted by Zipf’s Law [17]. Further, linguistic theory leads us to expect that the top-frequency words are function words (e.g., articles and prepositions), which is indeed what we observe here. However, looking at individual words in Fig. 6, we see aberrations. Two observations are important: First, for truncation masking the four most frequent words are markedly more frequent than with the other forms of masking. These are function words (“the”, “a”, and “of”) and one noun (“person”), which reflects the fact that sentences tend to start with these words. Second, for syntax masking certain nouns form peaks, meaning that they are more frequently retained with syntax masking than with other forms of masking.
Given the design of syntax masking, it is unsurprising that the words that are disproportionately retained belong to the POS category Noun. We dive deeper into the impact of existing masking approaches by examining the word distribution over POS categories.  Tab. 1 illustrates statistics for four major part-of-speech classes, Nouns, Adjectives, Verbs and Other (predominantly function words).
@@
Table 1: Part-of-speech (POS) distribution: The percentage of nouns (NN), adjectives ($J$), verbs (VB) and Other POS categories before and after text masking, calculated on CC12M, retaining 6 words per text.
We observe that the distribution for truncation, random, and block masking is comparable to the distribution before masking. However, for syntax masking, the number of words in the Noun category is extremely high and the number of words in the Other category is extremely low.
The prioritization of the POS category Noun and the deprioritization of Other has intuitive appeal. Although not explicitly mentioned, we assume that this appeal motivated CLIPA [20] in their design of syntax masking. Nouns are likely to be the name of an object directly depicted in an image, and can form a better bridge between image and text than words in the categories Adjectives or Verb. Further, words in the category of Other, being function words, are often classical stop words that are removed in information retrieval applications. Our results (Fig. 1) show that this intuition does not directly lead to good performance and must be revisited.
In short, our analysis reveals the importance of considering word frequency when text masking. We have seen that text masking approaches that do not disproportionately impact the frequency of certain words (i.e., random and block) outperform text masking strategies that do (i.e., truncation and syntax). The word frequency changes caused by using sentence-initial words (truncation) or emphasizing/de-emphasizing certain POS categories (syntax) hold back performance as the number of epochs increases. This observation motivates the importance of a text masking approach that is directly maintains relative word frequency, which we introduce in the next section.
4 Word Frequency Masking (CLIPF)
Contrastive Language-Image Pre-training with Word Frequency Masking (CLIPF) removes a word $w_{i}$ from the training data according to its masking probability score, $P(w_{i})$, which is calculated on the basis of the word’s frequency as:
$P(w_{i})=1-\sqrt{\frac{t}{f(w_{i})}}$ (3)
Word frequency, $f(w_{i})$, is the number of times the word appears in the training data divided by the total number of words in the training data. Eq. 3 scales $f(w_{i})$, resulting in a much higher masking probability for frequent words than infrequent ones. It also preserves the overall ranking of word frequencies.
The aggressiveness with which high frequency words are masked is controlled in Eq. 3 by the threshold $t$. CLIPF enjoys the benefit of being relatively insensitive to the parameter $t$, meaning that the parameter does not have to be precisely tuned. The influence of $P(w_{i})$ on $t$ is illustrated by a plot in the supplemental material, which also includes an experimental demonstration that the CLIPF is relatively insensitive to the choice of $t$.
The basic form of Eq. 3 has been validated previously in the literature, having been introduced by [29] for the purpose of text model training, which inspired CLIPF. Specifically, [29] and SW-CLIP [23] (which attempted to transfer [29] to CLIP training) discard words whose $P(w_{i})$ exceeds a cutoff. If there are not enough low $P(w_{i})$ words in a training text to fill in the input, these approaches use padding. CLIPF does away with this cutoff, and instead fills the input with words in order of $P(w_{i})$. Because CLIPF uses higher-$P(w_{i})$ words rather than padding it can outperform SW-CLIP, as shown in the supplementary material.
5 Performance of CLIPF
In this section, we discuss the performance of CLIPF on the basis of a reproduction of the CLIPA paper ( Sec. 5.1) with a model trained on LAION-400M [35] as well as the initial experiment in Fig. 1 with models trained on CC3M [36] and CC12M [3].
5.1 Revisiting CLIPA
We carry out a reproduction of the original CLIPA experiment on LAION-400M [35]. We remain as faithful as possible to the original setup, with two exceptions. First, not all 400M URLs from LAION-400M are still available online, and we train on the 298M pairs that are available. Second, we reduce the computational expense of the experiment by using a smaller image size (112x112) and 50% image masking. We employ a ViT-B/16 image encoder with global average pooling and learnable positional embeddings. As with CLIPA, we fine-tune for 0.4 epochs on the full data after training on the masked data. More details about the experimental setup are available in the supplementary material.
Results for zero-shot classification on ImageNet-1K are shown in Fig. 4 for both top-1 accuracy (bottom) and top-5 accuracy (top) for 4-16 epochs of training on the masked data. We observe that CLIPF outperforms syntax masking.
Figure 4: Zero-shot classification accuracy on ImageNet-1K using models trained on LAION-400M to which different text masking strategies have been applied. The backbone of the image encoder is ViT-B/16. We pre-trained the model using 25 image tokens and 4 text tokens for 16 epochs. At each point, we apply an additional 0.4 epoch of training on the full, unmasked data.
The original CLIPA paper established that truncation masking was not a good performer, and we do not test it here. However, we find that other aspects of the conclusions of CLIPA concerning text masking are limited. Specifically, in order to observe this difference it is necessary to continue training until the increase in performance starts to level off. The original CLIPA paper was trained with only 6.4 epochs, and here we train to 16. The additional epochs allow us to observe block masking and random masking moving towards syntax masking as the number of epochs increases, with block masking on par with syntax masking with respect to top-5 accuracy at 16 epochs.
Comparing Fig. 4 with the first part of Fig. 1, we see that the relative levels of performance of syntax, block, random and CLIPF are the same. This comparability gives us confidence that we can study text masking with smaller data sets, which allows us to train for a large number of epochs using different comparative conditions, and we carry out the remaining experiments on data sets that require less computational expenditure.
5.2 Advantages of CLIPF
We report on the performance of CLIPF vs. the original CLIP with an experiment carried out on CC3M [36] and CC12M [3]. Not all text-image pairs were available online and our final data sets consist of 2.7M pairs from CC3M and 9.3M pairs from CC12M. For CLIPF, we fine-tune for one epoch on the full data after training on the masked data. We employ the same ViT-B/16 image encoder, but we use a 224×224 image input size. Further details about the experimental setup are available in the supplementary material. The results in Tab. 2 confirm that CLIPF is able to outperform CLIP. “Image tokens” and “text tokens” reflect the amount of image and text processing during pre-training and this performance gain comes with the computational advantage of using half the number of image tokens and one quarter of the number of text tokens.
Models Masking Image Tokens Text Tokens CC3M CC12M
pre-train fine-tune pre-train fine-tune
CLIP 197 32 18.6 36.6
CLIPF frequency 98 8 19.6 18.7 39.8 41.0
Table 2: Comparison of zero-shot classification on ImageNet-1K between CLIP and CLIPF using 98 image tokens and 8 text tokens. Reported as top-1 accuracy. All models use a ViT-B/16 image encoder backbone. They are pre-trained on CC3M and CC12M for 30 epochs (pre-train). CLIPF is fine-tuned for an additional epoch on all (no masking) data (fine-tune).
When text masking in an application scenario, choices need to be made, as already mentioned in Sec. 1. As seen in Fig. 1, the number of epochs matters. Here, we choose to train with 30 epochs, following [15]. The aggressiveness of the masking used will be determined by the computational resources available in the application scenario. Here, we used 50% image masking and in our further experiments, we adopt 25% image masking for greater speed. In the next section, we carry out a quantified evaluation of all existing text masking approaches, comparing them to CLIPF. This evaluation provides the information necessary to support the choices of text masking approaches, including the number of text input tokens, for use in application scenarios.
6 Comparison of Text Masking Strategies
In this section, we carry out a systematic evaluation of the existing text masking approaches from FLIP [22] (random and truncation), CLIPA [20] (block and syntax) and our CLIPF approach (frequency-based). In a nutshell, the results support our observation that syntax and truncation masking are not productive approaches to masking and highlight the benefits of masking that directly leverages frequency information, i.e., our CLIPF. These experiments are carried out on CC3M [36] and CC12M [3] with the same experimental setup as in Sec. 5.2, using 30 epochs, as previously mentioned, following OpenCLIP [15]. Following prior evaluations of CLIP models [32, 22], we evaluate the models’ zero-shot performance on ImageNet-1k (Sec. 6.1), zero-shot robustness (Sec. 6.2), zero-shot image-text retrieval (Sec. 6.3) and zero-shot classification on 25 additional data sets (Sec. 6.4). Further details of the experimental setup are in the supplementary material.
@@ 6.1 Zero-shot Performance on ImageNet-1k
Tab. 3 presents results for zero-shot classification on ImageNet-1K, which complement the ImageNet-1K results already presented in Fig. 1. The experiments compare different masking approaches and different masking ratios, indicated in the table with the number of input text tokens (Text Tokens column). Models are pre-trained for 30 epochs (pre-train column) and then fine-tuned for one epoch (fine-tune column) with the full, unmasked data. Note that since the average text length in CC3M is less than 16 tokens, the benefits of fine-tuning a model pre-trained with 16 text tokens are limited.
From Tab. 3, we observe the ability of fine-tuned text-masking strategies to outperform FLIP [22] in many cases (image masking alone), while using fewer text input tokens. This result is important since [22] states that text masking does not contribute to improvement in model performance. We attribute the fact that we find a gain to the limited number of epochs (6.4) with which [22] tested text masking.
From Tab. 3, we also observe that CLIPF outperforms the other masking matching approaches in nearly all cases across text masking ratios (input sizes), as reflected by the bolded accuracy scores. Fine-tuned CLIPF delivers good performance both on CC3M and also the larger CC12M. On CC12M, fine-tuned CLIPF even outperforms the original CLIP (no masking) with a 75% masking ratio (8 input tokens). Note that at a text masking ratio of 75% (corresponding to 8 tokens), the masking approaches are 1.3 $\times$ faster than FLIP and 4.0 $\times$ faster than CLIP.
Truncation performs poorly, consistent with the observations of CLIPA [20]. However, in contrast to [20], we observe that syntax masking is a poor performer. In contrast, approaches that do not disproportionately impact the frequencies of words (random and block) perform more strongly.
@@
Finally, we consider the pattern of the text masking approaches as the masking ratio increases, i.e., as we decrease the input size from 16 to 4 tokens. Here, we focus on the results of the fine-tuned model. Moving from 16 to 8 input tokens benefits most approaches. The benefit derives from the fact that the reduction of the number of input tokens allows for an increase in the batch size. As we reduce the number of tokens from 8 to 6 to 4, the gap between fine-tuned CLIPF and the next best masking strategy widens. We see that compared to existing text matching approaches, CLIPF is better positioned to deliver benefits as the number of input tokens decreases.
Models Masking Image Tokens Text Tokens CC3M CC12M
pre-train fine-tune pre-train fine-tune
CLIP 197 32 18.6 36.6
FLIP 49 32 14.1 14.2 32.0 33.7
FLIP truncation 49 16 13.8 13.8 32.8 32.8
random 13.9 13.9 33.7 34.3
CLIPA block 13.9 13.9 34.3 34.8
syntax 13.3 12.8 32.2 34.4
CLIPF frequency 14.0 14.0 35.5 36.0
FLIP truncation 49 8 10.8 12.0 25.4 28.4
random 17.6 17.4 34.5 36.9
CLIPA block 16.2 16.6 35.5 37.9
syntax 17.2 17.5 28.5 35.0
CLIPF frequency 16.8 17.0 36.6 39.3
FLIP truncation 49 6 8.4 9.4 15.3 23.2
random 12.8 17.9 26.9 34.6
CLIPA block 12.9 17.0 28.6 35.9
syntax 12.2 15.7 25.2 32.6
CLIPF frequency 14.4 18.2 30.3 37.8
FLIP truncation 49 4 3.8 8.2 5.3 19.8
random 5.4 14.6 14.0 27.1
CLIPA block 7.5 14.5 18.7 26.6
syntax 8.9 13.0 14.2 24.6
CLIPF frequency 10.9 16.0 17.0 30.9
Table 3: Comparison of CLIP, FLIP, CLIPA and CLIPF for zero-shot classification on ImageNet-1K. Reported as top-1 accuracy. All models use a ViT-B/16 image encoder backbone pre-trained on CC3M and CC12M for 30 epochs with 75% image masking for speedup (pre-train) and fine-tuned for an additional epoch on all data (fine-tune). “Image tokens” and “text tokens” reflect the amount of image and text processing during pre-training.
@@ 6.2 Zero-shot Robustness Evaluation
Tab. 4 presents results for zero-shot classification for six data sets commonly used for robustness evaluation (e.g., by [32, 22]): ImageNet-A (IN-A) [13], ImageNet-O (IN-O) [14], ImageNet-R (IN-R) [12], ImageNet-Sketch (IN-S) [41], ImageNet-V2 (IN-Vs) [33], and ObjectNet (ON) [1]. The results lead to the same observations as with the zero-shot performance on ImageNet-1k in Sec. 6.1. Across different text masking ratios and across different data sets, CLIPF outperforms the other masking strategies in nearly all cases. Truncation and syntax are generally not good performers. Reducing the text tokens in CLIPF from 16 to 8 results in substantial performance improvements compared to the other text masking approaches across almost all datasets. Again, we see that the benefits of CLIPF are particularly evident as the masking number of input tokens decreases.
We also observe that CLIPF is able to improve over the FLIP baseline, without text masking in a majority of cases. However, despite being the strongest text-masking strategy, the results reveal that it is challenging for CLIPF to improve over the original CLIP. In only a handful of cases does it outperform the original CLIP, but might do so with more image tokens as we saw in Sec. 5.2.
Models Masking Image Tokens Text Tokens Text Retrieval Image Retrieval
@@
CLIP 197 32 62.62 35.54 45.42 24.22
FLIP 49 32 54.73 29.34 38.62 20.88
FLIP truncation 49 16 44.67 25.54 34.99 19.64
random 58.48 32.36 43.61 22.81
CLIPA block 56.51 30.82 44.06 22.66
syntax 54.54 29.60 41.16 21.40
CLIPF frequency 57.89 31.52 42.72 22.57
FLIP truncation 49 8 30.47 16.92 23.96 12.78
random 58.58 36.24 43.79 23.16
CLIPA block 60.06 35.88 45.98 24.65
syntax 50.30 29.64 38.46 20.28
CLIPF frequency 58.68 34.74 44.58 23.31
FLIP truncation 49 6 32.25 16.18 22.39 11.93
random 55.23 32.58 42.47 21.34
CLIPA block 54.44 33.12 41.68 21.73
syntax 46.15 26.78 34.08 17.99
CLIPF frequency 56.02 32.32 41.05 21.28
FLIP truncation 49 4 30.97 17.44 21.09 11.07
random 41.62 24.14 30.26 15.63
CLIPA block 40.83 23.56 30.77 14.98
syntax 38.66 21.68 26.08 14.14
CLIPF frequency 45.07 25.60 31.72 15.96
Table 5: Zero-shot Image-Text retrieval evaluation. We evaluated CLIP, FLIP, CLIPA, and CLIPF image-text retrieval performance on the COCO and Flickr30k datasets. The models are pre-trained on CC12M [3] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
@@
6.3 Zero-shot Image-Text Retrieval
Tab. 5 shows image-text retrieval results on Flickr30k [44] and COCO [24]. In general, we see that text masking approaches are able to improve over FLIP. When using 16 text tokens, all strategies except the truncation method outperform FLIP, which is pre-trained on the full text. Here, we do not see consistent improvement over CLIP, although some cases come close. Note again that we are testing here only with 49 image tokens and 98 would fare better (see Tab. 8 of the supplementary material), with additional computational cost.
We point out that for image-text retrieval the performance of syntax masking is particularly low in comparison to the other forms of masking. Specifically, when using 8 and 6 text tokens, the performance gap between syntax masking and block, random, and frequency masking is approximately 9% on COCO and 6% on Flickr30k for image-text retrieval. This gap is wider than what was observed with zero-shot classification. We conjecture that the bad performance is due to the image-text retrieval task requiring understanding the overall meaning and logic of the text rather than just recognizing nouns or objects. For this reason, the fact that syntax masking retains a high number of nouns, as shown in  Sec. 3, is particularly harmful.
6.4 Zero-shot Classification on Additional Datasets
In  Tab. 6, we present evaluation results on additional datasets. With six text tokens, CLIPF consistently outperforms the original CLIP (no masking), FLIP (image masking with full-text), and other text masking strategies. Aligned with ImageNet findings, reducing text tokens from 16 to 8 substantially boosts performance. Improvements include SUN397 [42] (+6.39%), DTD [6] (+9.31%), OxfordPets [31] (+4.64%), STL10 [7] (+9.31%), EuroSAT [11] (+6.60%), UCF101 [37] (+6.68%). The performance of other text masking strategies, including truncation, random, block, and syntax, also improves with this reduction. However, their improvements are more limited.
We point out that the Describable Textures Dataset (DTD)[6] contains 47 categories of texture patches, whose POS is mostly verb and that the OxfordPets dataset[31] also contains many class names that are not nouns. On these data sets, the performance of syntax masking is particularly low. Specifically, when using six text tokens, syntax masking underperforms CLIPF by 14.10% on DTD and 21.04% on OxfordPets. This low performance reflects that the VLM has overfitted on the large number of nouns retained by syntax masking and mistakes in POS detection (which is challenging for OxfordPets). Note that, unlike the performance on ImageNet (see Figs. 4 and 5 of the supplementary material), syntax masking on DTD remains worse than random, block, and frequency masking throughout all epochs.